OpenStudy (linn99123):

What is the slope of the graph of 6x + 12y = 7? A.) -2 B.) -1/2 C.) 1/2 C.) 2

OpenStudy (phebe):

-1/2

OpenStudy (anonymous):

You should at least show them how to find it @Phebe You have to get this into y=mx+b form so to do that you subtract 6x from both sides. That gives you 12y=-6x+7 Next divide both sides by 12 and you get y=-1/2x +7/12 Now referring to the formula y=mx+b, slope=m so the slope is -1/2

OpenStudy (phebe):

ok i will next time @Novalynne